The Fussy Food Snack Bar makes a trail mix of peanuts and raisins. If the cost of peanuts is $2.00 per pound and the cost of raisins is $3.50 per pound, how many pounds of peanuts and how many pounds of raisins should be used to make a 3 pound blend that costs $3.00 per pound?
Answer by palanisamy(496)

Let x pounds of peanuts and y pounds of raisins be used.
Total is x+y = 3 ...(1)
Also, the cost of peanuts is $2.00 per pound and the cost of raisins is $3.50 per pound.
The resulting blend costs $3 per pound
So, 2x+3.5y = 3*3
2x+3.5y = 9 ...(2)
(1)*2=> 2x+2y = 6 ...(3)
(2)-(3)=> 1.5y = 3
y = 3/(1.5)
y = 2
(1)=> x+2 = 3
x = 3-2
x = 1